OpenStudy (anonymous):

I rock is dropped of a cliff. 3.4 seconds later, you hear the sound it made hitting the ground. Let the speed of sound equal 340 m/s, and, neglecting air resistance, calculate the height of the cliff.

OpenStudy (anonymous):

What I tried was: \[\huge 3.4s = \frac {h}{340 \frac{m}{s}} + \sqrt{\frac {2h}{9.8\frac {m}{s^2}}}\] which I was able to simplify to: \[\huge 3618.85 = 3.13h + 340\sqrt{2y}\] I don't know how to solve it beyond that, but the answer, 52m, works in that equation.

OpenStudy (anonymous):

The h/340 is the portion of the time spent by the sound traveling up the cliff, and the sqrt(2h/9.8) is the time spent falling, taken from the equation \[\Delta y = \frac{1}{2}at ^{2}\]

OpenStudy (saifoo.khan):

Are u trying to impress us by making fractions?

OpenStudy (saifoo.khan):

lol, then keep in mind, we won't we getting impressed by this, Bring something new.

OpenStudy (anonymous):

That was a joke; I'm not trying to impress you with fractions. I'm trying to find out how to solve this problem

OpenStudy (saifoo.khan):

lol, but u said so.

OpenStudy (saifoo.khan):

-Just makin clear-

OpenStudy (anonymous):

So you don't know how to solve this, right?

OpenStudy (dumbcow):

ok i get 51.7, ......move the (h/340) to other side, then square both sides.....you will end up with a quadratic

OpenStudy (anonymous):

Ah, awesome, thanks dumbcow :)

OpenStudy (anonymous):

chris is awesome

OpenStudy (dumbcow):

btw, where did the y come from ??

OpenStudy (dumbcow):

thanks ishaan

OpenStudy (anonymous):

oops, that was supposed to be an h